# Integrated GARCH-GRU in Financial Volatility Forecasting

### v1

@@ 5 Applications
Formally, for a return series ${r_{t}}$, the one-period ahead VaR at significance level $\alpha\in(0,1)$ is defined as the quantile of the conditional return distribution such that the probability of a loss exceeding the VaR level is no greater than $\alpha$. Mathematically, the VaR at level $\alpha$, denoted $\text{VaR}_{t+1}^{\alpha}$, satisfies:
$\mathbb{P}(r_{t+1}<-\text{VaR}_{t+1}^{\alpha}\mid\mathcal{F}_{t})=\alpha$ (31)
where $\mathcal{F}_{t}$ represents the information set available at time $t$. Under the assumption that the return series have constant mean $\mu$, in this study, rather than assuming a parametric distribution (e.g., Gaussian), we estimate VaR using a non-parametric approach based on the empirical distribution of standardized returns. Specifically, we compute:
$\operatorname{VaR}_{t+1}^{\alpha}=\mu-q_{\alpha}(\varepsilon)\cdot\hat{\sigma}_{t+1}$ (32)
where $q_{\alpha}(\varepsilon)$ denotes the $\alpha$-quantile of the empirical distribution of $\varepsilon_{t+1}$, estimated from historical standardized residuals $\left\{\hat{\varepsilon}_{t}\right\}=\left\{\left(r_{t}-\mu\right)/\sigma_{t}\right\}$. And $\hat{\sigma}_{t+1}$ is the one-day-ahead forecasted volatility produced by the model. This method allows the VaR estimate to reflect the empirical distributional features of asset returns, such as skewness and fat tails, without imposing restrictive distributional assumptions.
To evaluate the quality of the VaR forecasts, we compute two diagnostic metrics: the violation count and the violation ratio. A violation occurs when the actual return falls below the estimated VaR threshold:
@@
indicating an underestimation of the true downside risk. The violation count is the total number of such instances, and the violation ratio is computed as:
@@
which is compared to the nominal level $\alpha$ to assess whether the model underestimates or overestimates risk.
Figure 3: Actual Returns, VaR Forecasts, and Violations at 95% Confidence Level on S&P 500
We evaluate and compare the 1-day-ahead VaR performance of seven volatility forecasting models: the proposed integrated GARCH-GRU and GARCH-LSTM models, the embedded hybrid variant (bl_GARCH_LSTM), two pipeline-based architectures (pl_GARCH_GRU and pl_GARCH_LSTM), and two standard econometric baselines: GARCH(1,1) and GJR-GARCH. All models are evaluated on the S&P 500 dataset. The actual returns and associated VaR forecasts are illustrated in Figure 3, and the violation statistics are summarized in Table 4.
Table 4: VaR Forecast Violations (One-Day Ahead Forecasts on S&P 500)
Model Violation Numbers Violation Ratio
eh_GARCH_GRU 3 1.30%
eh_GARCH_LSTM 8 3.48%
bl_GARCH_LSTM 15 6.52%
pl_GARCH_GRU 5 2.17%
pl_GARCH_LSTM 7 3.04%
GARCH(1,1) 18 7.83%
GJR-GARCH 17 7.39%
‘eh_GARCH_GRU’ and ‘eh_GARCH_LSTM’ denote the proposed end-to-end integrated models. ‘bl_GARCH_LSTM’ refers to the embedded integration approach that modifies the output gate within the LSTM architecture. ‘pl_GARCH_LSTM’ and ‘pl_GARCH_GRU’ represent pipeline-based integration models, where GARCH and neural components are trained sequentially.
Among all models, the integrated GARCH-GRU (eh_GARCH_GRU) model achieves the lowest violation count (3) and violation ratio (1.30%). In contrast, the baseline econometric models (GARCH(1,1) and GJR-GARCH) exhibit much higher violation ratios, 7.83% and 7.39%, respectively, indicating significant underestimation of downside risk. The other hybrid models also tend to generate more violations than the GARCH-GRU, suggesting less accurate tail risk modeling. These findings underscore the practical utility of the proposed GARCH-GRU model in risk-sensitive applications, particularly in capturing complex, nonlinear volatility dynamics and providing more reliable downside risk estimates.
@@ 6 Conclusion and Future Work
This study contributes to the advancement of financial volatility modeling by proposing a novel hybrid architecture that unifies traditional econometric structure with modern deep learning techniques. Specifically, we introduce an integrated GARCH-GRU model, which structurally embeds the GARCH(1,1) volatility dynamics directly into the recurrent computations of a GRU cell through a dedicated multiplicative gate. This architectural design enables the model to simultaneously capture key volatility stylized facts, such as clustering and persistence, while also learning complex nonlinear temporal dependencies inherent in financial time series.
Empirical evaluations conducted on major financial indices, S&P 500, Dow Jones Industrial Average, and NASDAQ, demonstrate the superior performance of the proposed framework. The integrated GARCH-GRU model consistently achieves lower forecasting errors (MSE and MAE) compared to a range of baseline models, including standalone neural networks (GRU, LSTM, Transformer), traditional GARCH-family models (GARCH(1,1), GJR-GARCH), and several recently proposed hybrid variants. The model not only offers improved statistical accuracy, but also shows computational efficiency, requiring less than 40% of the training time needed by its LSTM-based counterpart, making it particularly attractive for high-frequency or large-scale financial applications.
In addition to point forecast accuracy, we demonstrate the practical utility of the GARCH-GRU framework in risk management applications through Value-at-Risk (VaR) forecasting. The model produces the lowest violation ratio among all candidates in a one-day-ahead VaR backtesting experiment, aligning closely with the nominal risk level and outperforming both traditional and hybrid benchmarks. This result affirms the model’s ability to deliver reliable tail risk estimates and highlights its relevance for real-world deployment in regulatory or capital allocation settings.
Overall, the proposed integrated GARCH-GRU model demonstrates how a tightly coupled, end-to-end integration of econometric structure and neural network architecture can lead to substantial improvements in both forecasting accuracy and model interpretability. It advances beyond conventional hybrid approaches that treat econometric and machine learning components in isolation, offering a unified framework in which both are optimized jointly. This design enables the model to harness the complementary strengths of both paradigms, combining the statistical rigor of GARCH processes with the nonlinear modeling flexibility of deep learning, thereby yielding a more powerful and principled approach to volatility forecasting.
One current limitation of the proposed framework lies in the computational inefficiency introduced by the use of custom recurrent cell implementations. Although these customized cells have been functionally validated and the GARCH-GRU model demonstrates superior efficiency relative to the GARCH-LSTM, requiring significantly less training time, the reliance on non-native deep learning components incurs runtime overhead compared to standard, highly optimized neural modules. This gap becomes more pronounced in large-scale or low-latency applications. To address this, future work will focus on improving computational efficiency through code-level optimizations, operator fusion techniques, and hardware-aware implementation strategies, enabling the model to scale more effectively while retaining its modeling advantages.
Looking ahead, the flexibility of the proposed framework opens several promising directions for future research. One natural extension involves coupling the GRU architecture with asymmetric GARCH variants, such as EGARCH or GJR-GARCH, to better capture leverage effects and asymmetries in volatility responses to positive and negative shocks. Another avenue lies in the integration of attention mechanisms, which could enhance the model’s ability to focus on relevant historical patterns, particularly in high-frequency financial forecasting where data are noisy and regime-dependent. In addition, the framework could be extended to incorporate regime-switching dynamics, allowing it to model structural breaks and transitions between different volatility regimes over time. These enhancements would further deepen the integration of econometric rigor and machine learning flexibility, resulting in models that are not only highly interpretable and scalable, but also robust to the complexities of modern financial markets.
Appendix A In-Sample Analysis - GARCH Parameter Estimation Results
Table 5: In sample analysis - Parameters Estimation for 1-day ahead Forecast
Model eh_GARCH_GRU eh_GARCH_LSTM bl_GARCH_LSTM pl_GARCH_LSTM GARCH(1,1) GJR-GARCH
S&P 500 $\omega_{0}$ 0.7217 (5.450e-02) 0.6085 (3.490e-02) 0.3525 (1.495e-01) 0.0365 (8.248e-03) 0.0344 (8.969e-03) 0.0342 (6.468e-03)
$\alpha$ 0.3063 (9.290e-02) 0.0636 (1.070e-02) 0.0661 (5.810e-02) 0.1697 (2.513e-02) 0.1597 (2.644e-02) 0.0000 (3.487e-02)
$\beta$ 0.6837 (9.290e-02) 0.9264 (1.070e-02) 0.7610 (1.115e-01) 0.7914 (2.435e-02) 0.8051 (2.709e-02) 0.8270 (2.697e-02)
$\gamma$ 0.2642 (4.517e-02)
DJI $\omega_{0}$ 0.6920 (6.410e-02) 0.6416 (4.930e-02) 0.3485 (1.624e-01) 0.0342 (7.130e-03) 0.0316 (7.373e-03) 0.0317 (5.663e-03)
$\alpha$ 0.6283 (1.903e-01) 0.0505 (9.900e-03) 0.1108 (6.090e-02) 0.1708 (2.534e-02) 0.1656 (2.644e-02) 1.3416e-10 (2.789e-02)
$\beta$ 0.3617 (1.903e-01) 0.9395 (9.900e-03) 0.6821 (9.290e-02) 0.7892 (2.454e-02) 0.7998 (2.629e-02) 0.8275 (2.607e-02)
$\gamma$ 0.2576 (3.799e-02)
NASDAQ $\omega_{0}$ 0.7276 (3.400e-02) 0.5710 (5.560e-02) 0.4508 (1.915e-01) 0.0530 (1.150e-02) 0.0487 (1.229e-02) 0.0521 (9.666e-03)
$\alpha$ 0.6058 (7.090e-02) 0.0752 (3.120e-02) 0.0874 (5.420e-02) 0.1340 (2.075e-02) 0.1252 (2.141e-02) 0.0000 (4.129e-02)
$\beta$ 0.3842 (7.090e-02) 0.9148 (3.120e-02) 0.7615 (8.140e-02) 0.8189 (2.356e-02) 0.8327 (2.566e-02) 0.8372 (3.103e-02)
$\gamma$ 0.2210 (4.238e-02)

### v2

@@ 5 Applications
Formally, for a return series $\{r_{t}\}$, the one-step-ahead VaR at significance level $\alpha\in(0,1)$ is defined as the conditional quantile of the return distribution such that the probability of a loss exceeding this level is no greater than $\alpha$. We express $\operatorname{VaR}$ on the loss scale, so that $\operatorname{VaR}_{t+1}^{(\alpha)}>0$ represents the $(1-\alpha)$-quantile of the conditional loss distribution. Under this convention, a VaR exceedance occurs whenever the realized return falls below the negative of this threshold:
$\mathbb{P}\bigl(r_{t+1}<-\text{VaR}_{t+1}^{(\alpha)}\mid\mathcal{F}_{t}\bigr)=\alpha,$ (32)
where $\mathcal{F}_{t}$ denotes the information set at time $t$. Assuming a constant conditional mean $\mu$, we compute VaR by combining the model’s volatility forecast with a parametric quantile based on the standardized return distribution. In this study, we adopt a Student-t specification for standardized residuals, which better captures the heavy tails commonly observed in equity returns. Let the standardized residuals be
$\hat{\varepsilon}_{t}=\frac{r_{t}-\mu}{\sigma_{t}},$
and let $q_{\alpha}^{(t)}$ denote the $\alpha$-quantile of the fitted Student-t distribution. The one-step-ahead VaR is then given by
$\text{VaR}_{t+1}^{(\alpha)}=-\mu-q_{\alpha}^{(t)}\,\hat{\sigma}_{t+1},$ (33)
where $\hat{\sigma}_{t+1}$ is the volatility forecast obtained from the model.
To evaluate the quality of the VaR forecasts, we first employ coverage-based diagnostics: the violation count and the violation ratio. A violation occurs when the actual return falls below the estimated VaR threshold:
@@
indicating an underestimation of the true downside risk. The violation count is defined as:
 Violation Ratio = Number of Violations  Total Number of Forecasts $\text${ Violation Ratio $\frac$$\text${ Number of Violations $\text${ Total Number of Forecasts }} (35)
and is compared to the nominal level $\alpha$ to assess whether the model underestimates or overestimates risk.
In addition to coverage-based evaluation, we further assess the accuracy of the predicted quantiles using the Pinball loss (also known as quantile loss). Unlike the violation ratio, which measures calibration at a specific confidence level, the Pinball loss provides a strictly consistent scoring rule that evaluates both the sharpness and reliability of quantile forecasts. For a realized return $r_{t+1}$ and the predicted quantile $\widehat{\operatorname{VaR}}_{t+1}^{(\alpha)}$, the Pinball loss is defined as
@@
Lower values of $\mathcal{L}_{\alpha}$ indicate more accurate and better-calibrated quantile forecasts, making the Pinball loss a natural complement to traditional violation-based backtesting in our VaR evaluation.
Figure 4: Actual Returns & VaR Forecasts at 99% Confidence Level on S&P 500
Table 9: Volatility and Tail-Risk Evaluation
VaR Backtesting ($\alpha=0.01$)
Model Violation Rate Violations Pinball_loss
GARCH_GRU 0.0261 6 0.0296
GARCH_LSTM 0.0348 8 0.0326
bm_GARCH_LSTM 0.0261 6 0.0321
Transformer 0.0304 7 0.0286
GARCH 0.0261 6 0.0361
GJR-GARCH 0.0261 6 0.0330
The VaR backtest is based on 230 one-day-ahead forecasts over the 2019 test window.
Figure.4 visualizes the one-day-ahead 99% VaR forecasts from all models against the realized S&P 500 returns in 2019. A correctly calibrated model should generate a VaR envelope that consistently lies below (i.e., more negative than) the realized return series except on rare occasions corresponding to true tail events. Both GARCH-RNN models, particularly GARCH-GRU, produce VaR paths that closely and coherently track the depth and timing of downside risk. Their VaR curves widen promptly around the major loss episodes in August and October 2019, reflecting a rapid increase in predictive tail thickness as volatility rises. By contrast, the classical GARCH-type models show noticeably slower adjustments and tend to underestimate risk immediately before sharp drawdowns, resulting in several clustered exceedances. The pipeline hybrid (bm_GARCH_LSTM) exhibits noisier and less stable VaR paths, again indicating weaker joint modeling of volatility and tail behavior. The Transformer captures some local loss patterns but displays occasional under-reactions around abrupt volatility spikes, consistent with its limitations under nonstationary dynamics. Overall, the visual evidence confirms that embedding GARCH structure within the recurrent cell enables both GARCH-GRU and GARCH-LSTM to generate more coherent and timely tail-risk estimates.
Table.9 further quantifies tail-risk performance using standard VaR backtesting metrics at the 99% level. A well-calibrated model should exhibit a violation rate close to the nominal 1% level, a small number of exceedances, and low Pinball loss values that reflect sharp and well-aligned quantile forecasts. Among all models, GARCH-GRU delivers the most balanced and stable performance, with six violations and a violation rate of 0.0261, while maintaining tight Pinball loss scores. Its consistent behavior across both metrics indicates strong conditional tail calibration and reliable responsiveness to latent volatility dynamics. The GARCH-LSTM variant also performs competitively but shows slightly higher violation frequency and weaker quantile sharpness, suggesting that its additional gating depth does not always translate into superior tail alignment during moderate volatility regimes. The two-stage hybrid (bm_GARCH_LSTM) matches the violation rate of GARCH-GRU but suffers from noticeably higher Pinball loss, reinforcing the advantage of unified learning over sequential pipeline designs. The Transformer achieves the lowest Pinball loss among all models, indicating highly sharp quantile prediction. However, its higher violation count reveals underestimation of tail intensity, which is an imbalance commonly seen in purely data-driven architectures lacking explicit volatility structure. Classical GARCH and GJR-GARCH models attain acceptable violation frequencies but exhibit substantially higher Pinball losses, confirming limited flexibility by their parametric structure and reduced capacity to match the sharpness and adaptability achieved by the neural and hybrid architectures in quantile forecasting.
In short, these findings demonstrate that the GARCH-RNN models, especially GARCH-GRU, deliver superior tail-risk performance. Their VaR forecasts respond rapidly and consistently to changes in market volatility, producing quantiles that track downside risk with appropriate depth and timing. The backtesting results further confirm robust calibration and quantile sharpness. These properties make the proposed GARCH-RNN framework particularly suitable for practical risk-management tasks, including VaR/ES estimation, stress testing, and capital-adequacy assessment. Overall, the results thus affirm the model’s capacity to bridge the gap between statistical volatility modeling and actionable financial risk control.
@@ 6 Conclusion and Future Work
This study contributes to the advancement of financial volatility modeling by proposing a novel hybrid architecture that unifies traditional econometric structure with modern deep learning techniques. Specifically, we introduce two integrated variants, GARCH-GRU and GARCH-LSTM, in which GARCH(1,1) volatility dynamics are embedded directly into the recurrent computation via a dedicated multiplicative gate. This construction enables the models to capture volatility clustering, long memory, and nonlinear temporal dependence within a single unified trainable system. An important advantage of this design is that it preserves interpretable GARCH parameters ($\omega$, $\alpha$, $\beta$), allowing the models to retain clear economic meaning while benefiting from the representational flexibility of neural networks.
Comprehensive empirical evaluation across major U.S. equity indices (S&P 500, Dow Jones Industrial Average, NASDAQ) demonstrates that both GARCH–GRU and GARCH–LSTM consistently outperform a wide set of benchmarks, including classical GARCH-family models, pipeline-style hybrids, and neural baselines such as the Transformer. Across multiple risk and accuracy metrics (MSE, $R^{2}$, MAE, SMAPE), the GARCH-RNN models achieve substantially higher predictive accuracy and exhibit greater stability across forecasting horizons. Within the proposed family, the GARCH–GRU emerges as the primary and most versatile variant, offering the strongest performance in stable or moderately volatile environments and achieving the best overall accuracy–efficiency tradeoff. It requires less than two-thirds of the training time of the GARCH–LSTM due to its more compact gating structure. The GARCH–LSTM, while more computationally intensive, provides additional memory depth that becomes advantageous in highly turbulent periods, allowing it to match or surpass GARCH–GRU in certain stress-period settings. Nevertheless, the performance gap between these two variants is small compared with the considerable gains they jointly achieve over all benchmarks, confirming that the primary improvements arise from the unified embedding of GARCH dynamics within the recurrent architecture. Moreover, both models yield stable and economically plausible GARCH parameter estimates, confirming that the unified design retains interpretability and diagnostic transparency which is an advantage rarely available in purely data-driven deep learning models.
The risk-based evaluation further confirms the robustness of the proposed framework. In 99% VaR backtesting, the GARCH-RNN models deliver lower violation ratios and sharper quantile predictions than classical GARCH specifications. The GARCH-GRU exhibits particularly strong calibration, achieving violation frequencies close to the nominal level while maintaining competitive Pinball losses. These results underscore the framework’s ability to adapt to tail-risk behavior and its suitability for practical applications in risk management, stress testing, and capital adequacy assessment.
Overall, the GARCH-RNN framework demonstrates that a unified integration of econometric volatility structure and neural recurrent dynamics can yield models that are simultaneously interpretable, flexible, and empirically powerful. Unlike traditional hybrid approaches that combine econometric and machine-learning components in a sequential or loosely coupled manner, the proposed unified architecture allows both components to reinforce each other within a coherent modeling system. This combination of interpretability, stability, and nonlinear representational capacity positions the GARCH-RNN family as a bridge between classical financial modeling and modern machine learning.
A remaining limitation of the framework lies in the computational overhead associated with custom recurrent-cell implementations. Although the proposed models have been carefully validated and the demonstrate superior efficiency relative to benchmarks, custom operators remain slower than native deep-learning primitives in large-scale or latency-sensitive applications. Future work will focus on code-level optimization, operator fusion, and hardware-aware implementation strategies to improve runtime performance without altering the unified model structure.
Looking ahead, several promising avenues extend naturally from this work. Incorporating asymmetric GARCH components (EGARCH or GJR-GARCH) into the unified architecture may improve the modeling of leverage effects. Integrating attention mechanisms could further enhance pattern recognition, particularly in noisy or high-frequency environments. Regime-switching or state-space extensions would allow the model to adapt to structural breaks and evolving volatility regimes. These directions hold the potential to enrich the unified econometric–neural design, yielding models that are not only interpretable and scalable but also robust to the increasingly complex dynamics of modern financial markets.
Data availability statement:
The data that support the findings of this study are openly available at Yahoo Finance.
Declaration of interests statement:
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix - Further Experimental Results
Table 10: Forecasting Performance in Normal-Period (Metric: SMAPE)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1 0.1653 (0.0060) 0.1603 (0.0061) 0.2771 (0.0066) 0.1529 (0.0064) 0.5394 0.5623
3 0.2206 (0.0105) 0.2236 (0.0095) 0.3194 (0.0080) 0.2336 (0.0056) 0.3868 0.3738
7 0.3056 (0.0090) 0.3149 (0.0104) 0.3815 (0.0203) 0.3058 (0.0074) 0.4559 0.4642
DJI 1 0.1803 (0.0150) 0.1617 (0.0078) 0.2806 (0.0120) 0.1688 (0.0142) 0.5097 0.5405
3 0.2311 (0.0124) 0.2235 (0.0066) 0.3075 (0.0007) 0.2481 (0.0078) 0.3621 0.3754
7 0.3169 (0.0056) 0.3202 (0.0073) 0.3603 (0.0060) 0.3088 (0.0121) 0.4203 0.4489
NASDAQ 1 0.1734 (0.0148) 0.1591 (0.0041) 0.2810 (0.0106) 0.1499 (0.0058) 0.5400 0.5608
3 0.2249 (0.0127) 0.2269 (0.0074) 0.3253 (0.0060) 0.2270 (0.0036) 0.3875 0.3799
7 0.3122 (0.0066) 0.3148 (0.0092) 0.3735 (0.0055) 0.3102 (0.0044) 0.4642 0.4583
Results are based on the 2019 out-of-sample evaluation. Reported values denote the mean and standard deviation across 20 seeds where applicable.
Table 11: Forecasting Performance in Normal-Period (Metric: MAE)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1 0.0945 (0.0035) 0.0963 (0.0044) 0.1757 (0.0040) 0.0890 (0.0038) 0.2174 0.2124
3 0.1343 (0.0059) 0.1389 (0.0068) 0.2042 (0.0051) 0.1146 (0.0037) 0.2529 0.2519
7 0.1989 (0.0062) 0.2071 (0.0084) 0.2558 (0.0176) 0.2000 (0.0042) 0.3125 0.3322
DJI 1 0.1073 (0.0092) 0.0954 (0.0047) 0.1780 (0.0068) 0.0991 (0.0095) 0.2101 0.2250
3 0.1423 (0.0078) 0.1390 (0.0052) 0.1970 (0.0042) 0.1536 (0.0050) 0.2396 0.2601
7 0.2098 (0.0050) 0.2125 (0.0064) 0.2393 (0.0053) 0.2033 (0.0093) 0.2906 0.3257
NASDAQ 1 0.1288 (0.0125) 0.1190 (0.0041) 0.2219 (0.0075) 0.1137 (0.0046) 0.2761 0.2766
3 0.1732 (0.0108) 0.1786 (0.0070) 0.2268 (0.0075) 0.1785 (0.0030) 0.3213 0.3233
7 0.2552 (0.0064) 0.2587 (0.0093) 0.3098 (0.0048) 0.2567 (0.0045) 0.4017 0.4169
Results are based on the 2019 out-of-sample evaluation. Reported values denote the mean and standard deviation across 20 seeds where applicable.